OpenStudy (eyebrowsfleek):

which expression represents the series 1+5+25+125+625

OpenStudy (anonymous):

Each term is 5 raised to a different power. 5^0 = 1 5^1 = 5 5^2 = 25 5^3 = 125 Ad infinitum.

OpenStudy (taylor0402):

The series 1+5+25+125+625 can be represented by the expression f(x) = 5^(x - 1).

OpenStudy (anonymous):

@Taylor0402 You don't need the x-1. You just need 5^x.

OpenStudy (eyebrowsfleek):

also thank you @Taylor0402